Help.....I am having a hard time in my finance management class and our book is not the best help

Lear, Inc., has $880,000 in current assets, $390,000 of which are considered permanent current assets. In addition, the firm has $680,000 invested in fixedassets.

(a)

Lear wishes to finance all fixed assets and half of its permanent current assets with long-term financing costing 9 percent. The balance will be financed withshort-term financing, which currently costs 7 percent. Lear’s earnings before interest and taxes are $280,000. Determine Lear’s earnings after taxes under thisfinancing plan. The tax rate is 30 percent. (Omit the "$" sign in your response.)

Earnings after taxes $

(b)

As an alternative, Lear might wish to finance all of its fixed assets and permanent current assets plus half of its temporary current assets with long-term financingand the balance with short-term financing. The same interest rates apply as in part a. Earnings before interest and taxes will be $280,000. What will be Lear’searnings after taxes? The tax rate is 30 percent. (Omit the "$" sign in your response.)

Earnings after taxes $

I would like to see the steps involved in solving this problem. Everything I have done is wrong and I don't understand why it is wrong. The textbook gives no examplesor guidance just very generic breakdowns (not sure why my school is using this book).

Thanks
Answer #1
sorry above one is wrong value solution. this one is correct.
all the negative values are in brackets
a)
Fixed Asset = $680,000
Permanent Current Asset= $390,000
Non-Permanent current Asset= $490,000

Total Asset financed through long term debt
= $680,000+$390,000*50% = $875,000

Interest on Long term debt= $875,000*9%= $78,750

Total Asset financed through short term debt
=$490,000+390,000*50%= $685,000

Interest on short term debt= $685,000*7%=$47,950

Earning Before Interest and Tax.....................280,000
Less: Interest on Long term debt.....................(78,750)
Less: Interest on Short term debt....................(47,950)
Profit Before Tax...........................................153 300
Less: Taxes..................................................(45 990)
Profit after tax..............................................107 310

b) Fixed Asset = $680,000
Permanent Current Asset= $390,000
Non-Permanent current Asset= $490,000

Total Asset financed through long term debt = $680,000+390,000+490000*50% = $1,315,000
Interest on Long term debt= $1,315,000*9%= $118,350

Total Asset financed through short term debt =$490,000*50%= $245,000
Interest on short term debt= $245,000*7%=$17,150

Earning Before Interest and Tax.....................$200,000
Less: Interest on Long term debt.....................(118,350)
Less: Interest on Short term debt....................(17,150)
Profit Before Tax...........................................64,500
Less: Taxes..................................................(19,350)
Profit after tax..............................................45,150
answered by: David Earnest
Answer #2
a)
Fixed Asset = $600,000
Permanent Current Asset= $350,000
Non-Permanent current Asset= $450,000

Total Asset financed through long term debt
= $600,000+350,000*50% = $775,000

Interest on Long term debt= $775,000*10%= $77,500

Total Asset financed through short term debt
=$450,000+350,000*50%= $625,000

Interest on short term debt= $625,000*5%=$31,250

Earning Before Interest and Tax.....................$200,000
Less: Interest on Long term debt.....................(77,500)
Less: Interest on Short term debt....................($31,250)
Profit Before Tax...........................................$91,250
Less: Taxes..................................................($27375)
Profit after tax..............................................$63,875

b) Fixed Asset = $600,000
Permanent Current Asset= $350,000
Non-Permanent current Asset= $450,000

Total Asset financed through long term debt = $600,000+350,000+450000*50% = $1,175,000
Interest on Long term debt= $1,175,000*10%= $117,500

Total Asset financed through short term debt =$450,000*50%= $225,000
Interest on short term debt= $225,000*5%=$11,250

Earning Before Interest and Tax.....................$200,000
Less: Interest on Long term debt.....................(117,500)
Less: Interest on Short term debt....................($11,250)
Profit Before Tax...........................................$71,250
Less: Taxes..................................................($21,375)
Profit after tax..............................................$49,875
answered by: nono